10. 1/3 points1 Previous Answers Priviter a Stats3 12 E.033 Using a rating scale, Tekinarslan (2008) measured computer anxiety among university students who use the computer very often, often, sometimes, and seldom. Below are the results of the one-way ANOVA Va 653.20 21.16 Between groups Within groups (errr 3,14.6110230.96 Total 1,959.79 5,108.41 105 a) What are the values for N and k? N- b) Was the sample size, ,eq each group? xplan. Yes, the sample size in each group was equal because such equality is necessary for one-way ANOVA O Yes, the sample size in each group was equal because the total sample size can be equally divided into the correct number of groups O No, the sample size in each group was nt equal because such inequality is necessary far one-way ANOVA No, the sample size in each group was not equal because the total sampla size can not be equally divided into the correct number of groups. Additional Materials eBook -1 points PriiteraStats3 13E005. which source of error is includad in the F-table for the one-way within-subjects ANOVA hut no for the one-way betwean-subjacts ANOVA? between- groups varlation within persons variation Additional Materials
Answer #1

10)

A)
N = 105 +1 = 106
K = 3+1 = 4

b)
option D) is correct
as 106 is not divided by 4

11)

d) Between persons variation is not included for between subjects ANOVA.